# MedThink: Explaining Medical Visual Question Answering via Multimodal Decision-Making Rationale

### v1

@@ 4.1 Setting
R-SLAKE Explanation 53.1 22.7 31.7 39.2 24.1 15.4 9.9
Reasoning 53.5 22.8 32.1 39.5 24.3 15.5 10.0
@@
Table 2: Accuracy (%) Comparison of Methods on Closed-End Questions in the R-RAD and R-SLAKE Datasets. Gray and red backgrounds indicate results for established and our methods.
{tblr}
cells = c, colspec = ccc, row1 = font=, row2-8 = bg=gray!10, row9-12 = bg=red!10, hline1,13 = 1pt,solid, vline2-3 = solid, Methods R-RAD R-SLAKE
MFB [?] 74.3 75.0
SAN [?] 69.5 79.1
BAN [?] 72.1 79.1
MEVF+SAN [?] 73.9 78.4
MEVF+BAN [?] 77.2 79.8
MMBERT [?] 77.9 -
PubMedCLIP [?] 79.5 82.5
w/o R 79.0 82.5
Reasoning 73.9 80.8
Two-Stage Reasoning 80.5 79.1
Explanation 83.5 86.3
In our framework, the encoder and decoder from UnifiedQA [?] are integrated as TextualEncoder($\cdot$) and TextualDecoder($\cdot$), respectively. Additionally, DETR [?] is employed as VisualEncoder($\cdot$). This integration facilitates the construction of our experimental model. Our experimental procedure is systematically divided into distinct phases. We initially evaluate the performance of the “Explanation”, “Reasoning” and “Two-Stage Reasoning” methods on the R-RAD and R-SLAKE datasets, benchmarking them against established models like PubMedCLIP [?], MMBERT [?], MEVF [?], BAN [?], SAN [?], MFB [?]. Subsequently, the quality of the rationales generated through our proposed methods, was assessed.
In our experiments, the learning rate is uniformly set at 5e-4 for both the R-SLAKE and R-RAD datasets. However, the number of epochs in the fine-tuning phase differs, comprising 300 epochs for the R-SLAKE dataset, but only 150 epochs for the R-RAD dataset. It is important to note that our “Two-Stage Reasoning” method requires a phased fine-tuning process involving two separate models. For the first model, we follow the parameters mentioned earlier. For the second model, we consistently use a learning rate of 5e-5, fine-tuning it for 20 epochs for both datasets. All experiments reported in this paper are conducted using PyTorch on an Ubuntu server equipped with eight NVIDIA RTX 3090 GPUs. The code and datasets from our study will be released.
4.2 Main Results
Our performance evaluation is divided into two parts, focusing separately on closed-end and open-end questions. Closed-end questions, structured as multiple-choice questions with a single definitive answer, are assessed using accuracy as the performance metric, as shown in Table 4.1. In contrast, open-end questions allow for a range of answers due to their inherent nature. The answers generated by our methods are difficult to match precisely against the dataset. Therefore, we employ text generation metrics like Rouge and BLEU to evaluate our methods’ performance, as exhibited in Table 4.1.
In facing closed-end questions, the “Explanation” method surpasses the “Reasoning” and “Two-Stage Reasoning” methods with accuracy of 83.5% on the R-RAD dataset and 86.3% on the R-SLAKE dataset, marking improvements of 4.0% and 3.8%, respectively, over the state-of-the-art PubMedCLIP model.
With open-end questions, different methods show distinct advantages. The Rouge scores, similar to the “Recall”, emphasizes the completeness of the generated text, while the BLEU scores, akin to the “Precision”, stresses the preciseness of the generated text. The “Explanation” method demonstrates higher Rouge and BLEU scores on the R-RAD dataset, with Rouge-1, Rouge-L, BLEU-1, BLEU-2, and BLEU-3 reaching 50.2%, 29.5%, 38.3%, 22.9%, and 14.0%, respectively. The “Two-Stage Reasoning” method showcases higher scores on the R-SLAKE dataset, with Rouge-2, BLEU-1, BLEU-2, BLEU-3, and BLEU-4 at 23.1%, 23.1%, 39.5%, 24.5%, 15.8%, and 10.3%, respectively. The “Reasoning” method maintains robust performance across both the R-RAD and R-SLAKE datasets; on the R-RAD dataset, Rouge-2, BLEU-3, and BLEU-4 reached 20.3%, 14.0%, and 8.9%, respectively, while on the R-SLAKE dataset, Rouge-1, Rouge-L, and BLEU-1 are 53.5%, 32.1%, and 39.5%, respectively. These outcomes highlight the necessity of diverse methods for various types of open-end questions.
4.3 Rationale Quality Assessment
@@
To further compare the “Explanation”, “Reasoning”, and “Two-Stage Reasoning” methods, we design two sets of experiments for closed-end questions on the R-RAD and R-SLAKE datasets. In the first set, we introduce a control experiment “w/o R”, where the models are trained and inferred without incorporating medical decision-making rationales. The “Explanation”, “Reasoning” and “Two-Stage Reasoning” methods are compared with the control experiment. As indicated in Table 4.1, relative to “w/o R”, the “Explanation”, “Reasoning”, and “Two-Stage Reasoning” methods showed accuracy improvements of 4.5%, -5.1%, and 1.5% respectively on the R-RAD dataset, and 3.8%, -1.7%, and -3.4% respectively on the R-SLAKE dataset.
In the second experimental series, we employ Gemini Pro to evaluate the assistance of medical decision-making rationales generated through our methods. Initially, the input of Gemini Pro is restricted to medical queries and associated imagery. Subsequently, medical decision-making rationales generated by the “Explanation”, “Reasoning” and “Two-Stage Reasoning” methods are incorporated to assist Gemini Pro in answering. We apply the self-consistency protocol [?], where Gemini Pro answers each question five times, selecting the most frequent answer as the final answer. The results, presented in Table LABEL:tab:_gemini_closed-end_acc, indicate an initial accuracy of 73.2% on the R-RAD dataset and 72.8% on the R-SLAKE dataset for Gemini Pro. The integration of medical decision-making rationales has led to significant improvements. Among three methods, the “Explanation” method stands out, enhancing the accuracy by 8.8% on the R-RAD dataset and 8.5% on the R-SLAKE dataset.
4.4 Ablation Study
To validate the impact of MLLM selection and expert annotation involvement on data quality during the data annotation process, we design three variations for annotating the closed-end questions in the R-RAD dataset: using Gemini Pro without expert input, using GPT-4V without expert input, and using GPT-4V with expert input. The results, as depicted in Figure LABEL:fig:_expert, indicate that GPT-4V enhances MedVQA task performance over Gemini Pro, aligning with prior researches [??] that illustrate GPT-4V’s more advanced reasoning abilities. Additionally, the participation of expert annotations is found to improve data quality.
4.5 Case Study
To observe the assistance of medical decision-making rationales in MedVQA tasks specifically, Figure LABEL:fig:_case_study shows several examples where the model employs the “Explanation” method. For instance, when addressing the question “Are the small bubbles of air seen abnormal?”, the model without Medical Decision-Making Rationales incorrectly responds. In contrast, the model utilizing the “Explanation” method not only provides the correct answer but also supplies the requisite medical knowledge, stating “The small bubbles appear as very dark or black on the CT scan”. Furthermore, it delivers pertinent descriptions of the medical image, noting “there are multiple areas of radiolucency, which are consistent with the appearance of small bubbles of air”. This underscores the supportive role of medical decision-making rationales in guiding the model to answer MedVQA questions accurately.
5 Conclusion
@@
In this paper, we present a generative model-based framework for MedVQA and construct the R-RAD and R-SLAKE datasets, which include intermediate reasoning steps to address the challenge of black-box decision-making processes in MedVQA models. Extensive experimental results show that our proposed framework not only elucidates the medical decision-making process of MedVQA models with clarity but also significantly enhances their performance. Future research will further explore generative models tailored for real clinical settings and how to better evaluate the performance of MedVQA models in open-ended scenarios.

### v2

@@ V-A Training Details
Explanation 50.2 20.2 29.5 38.3 22.9 14.0 8.8
R-SLAKE Reasoning 53.5 22.8 32.1 39.5 24.3 15.5 10.0
@@
Explanation 53.1 22.7 31.7 39.2 24.1 15.4 9.9
R-Path Reasoning 41.5 13.0 24.8 31.8 17.0 9.6 5.7
Two-Stage Reasoning 41.7 13.2 24.9 32.1 17.1 9.7 5.8
Explanation 41.9 13.2 25.0 32.1 17.1 9.7 5.8
All experiments reported in this paper are conducted using PyTorch on an Ubuntu server equipped with four NVIDIA RTX 3090 GPUs. Training on the R-RAD dataset takes about 2.5 hours. In comparison, training on the R-SLAKE dataset requires approximately 5.5 hours, while the R-Path dataset takes around 14 hours. During inference, processing each sample takes about 6 seconds.
V-B Evaluation Metrics
Our performance evaluation is divided into two parts, focusing on closed-end and open-end questions separately. For closed-end questions, which are formatted as multiple-choice with a single correct answer, we assess performance using accuracy as the metric. For open-end questions, in contrast to previous works [29, 30, 28, 10, 31, 13] that often emphasize scoring all possible answers in open-ended MedVQA datasets to gauge classification accuracy, our work on generative MedVQA prioritizes clinical utility. Following established studies [38, 39], we employ BLEU and ROUGE to assess the quality of our method’s outputs. The BLEU scores, akin to the “Precision”, evaluates the overlap of k-grams between generated and reference sentences, while the ROUGE scores, similar to the “Recall”, measures the similarity in word sequences.
V-C Main Results
In facing closed-end questions, we evaluate the performance of MedThink under various generation strategies, and compare them against several baseline methods on the R-RAD, R-SLAKE, and R-Path datasets. The results are shown in Table II. MedThink demonstrates varying levels of performance across different generation strategies. To be specific, MedThink with the “Explanation” strategy achieves the highest accuracy on the R-RAD and R-SLAKE datasets, recording 83.5% and 86.3% respectively. Meanwhile, MedThink with the “Two-Stage Reasoning” strategy achieves the best performance on R-Path with an accuracy of 87.2%.
In contrast, the state-of-the-art classification model, PubMedCLIP, achieves accuracies of 79.5% on the R-RAD dataset and 82.5% on the R-SLAKE dataset, which is significantly lower than MedThink’s results. This underscores the superior performance of MedThink. Compared to other generative models based on MLLM, MedThink outperforms models such as LLaVA [27], Med-Flamingo [32], and Med-Gemini [33], achieving overall accuracies that are on par with the more parameter-heavy LLaVA-Med [27] and Med-MoE [26] models. Notably, MedThink accomplishes this with a parameter count that is less than one-tenth of these models, demonstrating both its efficiency and effectiveness.
Using open-end questions, we conduct a comprehensive evaluation of MedThink’s three strategies on the R-RAD, R-SLAKE, and R-Path datasets. The results are summarized in Table III. For the R-RAD dataset, the ”Explanation” strategy outperforms other strategies, achieving the highest scores in five out of seven metrics. It records 50.2% in Rouge-1, 29.5% in Rouge-L, 38.3% in BLEU-1, 22.9% in BLEU-2, and 14.0% in BLEU-3. On the R-SLAKE dataset, the ”Two-Stage Reasoning” strategy leads in performance, securing the top scores in five out of seven metrics, with 23.1% in Rouge-2, 39.5% in BLEU-1, 24.5% in BLEU-2, 15.8% in BLEU-3, and 10.3% in BLEU-4. Regarding the R-Path dataset, the ”Explanation” strategy once again delivers the highest overall performance, achieving 41.9% in Rouge-1, 13.2% in Rouge-2, 25.0% in Rouge-L, 32.1% in BLEU-1, 17.1% in BLEU-2, 9.7% in BLEU-3, and 5.8% in BLEU-4. These results collectively highlight the importance of selecting appropriate generation strategies tailored to addressing various medical scenarios, ensuring the model generates comprehensive and detailed responses.
V-D Ablation Study
To explore the influence of various components in MedThink, we conduct a series of ablation experiments. First, we evaluate the effect of different MLLMs used during dataset creation and the contribution of domain experts to data annotation. We implement three variations for annotating the closed-end questions in the R-RAD dataset: using Gemini Pro [17] without expert involvement, using GPT-4V without expert involvement, and using GPT-4V with expert involvement. As shown in Figure LABEL:fig:_expert, when GPT-4V is used with expert involvement in dataset creation, the “Explanation” and “Two-Stage Reasoning” strategies achieved their highest accuracies of 83.5% and 80.5%, respectively. In contrast, the ’Reasoning’ strategy performed best with Gemini Pro without expert involvement, reaching an accuracy of 79.4%, which is only slightly above the baseline accuracy of 79.0% when MDMRs are not applied. We attribute this to the instability of the ’Reasoning’ strategy, which hinders its ability to consistently benefit from MDMRs, aligning with previous research [7]. Overall, expert involvement enhances the quality of MDMRs, positively impacting MedThink. Additionally, GPT-4V’s stronger reasoning ability compared to Gemini Pro [40] further suggests that using a more advanced MLLM during data annotation is beneficial.
@@ V-D Ablation Study
Next, we examine how the introducing of MDMRs impacted results. We introduce a control experiment, MedThink without MDMRs, where the models are trained and inferred without incorporating MDMRs (MedThink w/o R). The “Explanation”, “Reasoning” and “Two-Stage Reasoning” strategies are compared with the control experiment. As indicated in Table II, compared to “MedThink w/o R”, the “Explanation”, “Two-Stage Reasoning”, and “Reasoning” strategies improve accuracy by 4.5%, 1.5%, and -5.1% on the R-RAD dataset, 3.8%, -3.4%, and -1.7% on the R-SLAKE dataset, and 1.0%, 1.2% and -2.9% on the R-Path dataset, respectively.
Finally, we assess the practicality of MDMRs generated by MedThink using different strategies. Initially, Gemini Pro is provided with only medical queries and related imagery. Subsequently, we inincorporate MDMRs generated by MedThink with the “Explanation”, “Reasoning” and “Two-Stage Reasoning” strategies to assist Gemini Pro in answering. The results, presented in Table II, indicate an initial accuracy of 73.5% on the R-RAD dataset, 69.0% on the R-SLAKE dataset and 64.8% on the R-Path dataset for Gemini Pro. The integration of MDMRs has led to significant improvements. Among three strategies, the “Explanation” strategy stands out, enhancing the accuracy by 6.3% on the R-RAD dataset, 9.1% on the R-SLAKE dataset and 7.8% on the R-Path dataset.
V-E Case Study
To assess the specific impact of MDMRs on the MedVQA task, Figure LABEL:fig:_case_study shows several examples where MedThink applies the “Explanation” strategy to answer questions from the R-SLAKE datasets.
When the generated MDMR is accurate, MedThink can effectively and precisely answer the related medical question. If the MDMR contains errors, however, it misguides MedThink, leading to a phenomenon known as hallucination, which is a common issue in vision-language models. To investigate the causes of hallucinations in MedThink, we analyze the number of incorrect answers it provided on the R-SLAKE dataset. The R-SLAKE dataset is chosen because it covers medical questions about six anatomical regions, offering a complex and representative challenge.
We perform the analysis through the following steps. First, we categorize the test set questions by the anatomical regions associated with the medical images. Next, we tally the number of incorrectly predicted questions for each anatomical region. Finally, we calculate the proportion of incorrect predictions for each region, as shown in Table IV. The results indicate that MedThink significantly aids in addressing medical issues related to the chest and abdomen. However, these areas still account for the majority of prediction errors. We attribute this to the greater complexity of chest and abdomen images, which contain more organs than other regions, presenting a considerable challenge for the model.
VI Conclusion
@@
In this paper, we present a generative model-based framework for MedVQA and construct the R-RAD, R-SLAKE and R-Path datasets, which include intermediate reasoning steps to address the challenge of black-box decision-making processes in MedVQA models. Extensive experimental results show that our proposed framework not only elucidates the medical decision-making process of MedVQA models with clarity but also significantly enhances their performance. Future research will further explore generative models tailored for real clinical settings and how to better evaluate the performance of MedVQA models in open-ended scenarios.
TABLE IV: The error rate for each region (Lower values are better)
Anatomical Regions (Number) w/o Rationale $\downarrow$ Explanation $\downarrow$
Lung (N=141) 12.06% 9.93%
Abdomen (N=141) 24.11% 19.15%
Head (N=91) 18.68% 13.18%
Neck (N=16) 18.75% 6.25%
Chest (N=5) 20.00% 0.00%
Pelvic Cavity (N=22) 4.55% 13.64%